# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v1

@@ 4.1 Energy market model
$k=1,\dots,n$, where $p$ is the period, and $b_{0k},b_{1k},b_{2k},b_{3k}\in\mathbb{R}$. Let ${\bf X}\sim OU\text{-}VGC^{n}_{m}(\lambda,{\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ with initial value ${\bf X}_{0}$ in its stationary distribution. Thus, the BDLP is ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, and we work within this more general setting throughout, which includes ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a subclass. In Section 5.1, we make an additional assumption on $\boldsymbol{\eta}$, which is not needed now.
@@ 4.2 Esscher transform
It is immediate that $\mathbb{Q}_{{\bf h}}$ exists if and only if ${\bf h}\in{\cal D}_{\bf Z}$ and it is equivalent to $\mathbb{P}$. The Esscher transform is structure preserving in the sense that ${\bf Z}$ is also a Lévy process under $\mathbb{Q}_{\bf h}$, where ${\bf h}$ is commonly called the market price of risk (MPR).
The no arbitrage condition that the discounted traded asset prices are $\mathbb{Q}_{{\bf h}}$-martingales trivially holds for the risk-free asset, which is the only traded asset, so that the Esssher transform provides up to infinitely many EMMs $\mathbb{Q}_{{\bf h}}$ indexed by ${\bf h}\in{\cal D}_{\bf Z}$, all consistent with no arbitrage. The market chooses a MRP ${\bf h}\in{\cal D}_{{\bf Z}}$ based on risk preferences, which we infer by calibrating the model to observed forward and call prices.
@@
Without loss of generality, suppose $\boldsymbol{\eta}={\bf 0}$. Let ${\bf h}\in{\cal D}_{{\bf Z}}$. Combining (shir, , Section VII.2c, Theorem 1), that ${\bf V}_{1},\dots,{\bf V}_{m}$ are independent under $\mathbb{P}$, and Kac’s theorem, it follows that ${\bf V}_{1},\dots,{\bf V}_{m}$ are independent under $\mathbb{Q}_{{\bf h}}$. Specifically, with $\widetilde{\bf V}:=({\bf V}_{1},\dots,{\bf V}_{m})$, this claim follows from
$\kappa^{{\mathbb{Q}}_{\bf h}}_{\widetilde{\bf V}}(\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{m})$ $=\kappa_{\widetilde{\bf V}}(\boldsymbol{\theta}_{1}+{\bf h},\dots,\boldsymbol{\theta}_{m}+{\bf h})-\kappa_{\widetilde{\bf V}}({\bf h},\dots,{\bf h})$
$=\sum_{j=1}^{m}\kappa_{{\bf V}_{j}}(\boldsymbol{\theta}_{j}+{\bf h})-\kappa_{{\bf V}_{j}}({\bf h})$
$=\sum_{j=1}^{m}\kappa_{{\bf V}_{j}}^{\mathbb{Q}_{{\bf h}}}(\boldsymbol{\theta}_{j}),$
where as in the proof of Lemma 4.1, there exists a neighborhood $U_{j}\subseteq\mathbb{R}^{n}$ about ${\bf 0}$, such that $\boldsymbol{\theta}_{j}\in U_{j}$ and $\boldsymbol{\theta}_{j}+{\bf h}\in{\cal D}_{{\bf V}_{j}}$ for all $j=1,\dots,m$. Also, note that (4.4) holds since $\mathbb{E}[e^{\langle(\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{m}),\widetilde{\bf V}\rangle}]=\prod_{j=1}^{m}\mathbb{E}[e^{\langle\boldsymbol{\theta}_{j},{\bf V}_{j}\rangle}]$ with the LHS finite if and only if each term in the product is, otherwise both sides are infinite.
Thus, applying Lemma 4.1 to each ${\bf V}_{j}$ completes the proof. ∎
Note that all but the last line of this proof shows the fact that the Esscher transform preserves independence and is applicable to any Lévy processes.
@@
One way to see this is to take the example $n=2$, $a=0.5$, $\boldsymbol{\alpha}=(1,1)$, $\boldsymbol{\mu}=\boldsymbol{\eta}={\bf 0}$, ${\boldsymbol{\Sigma}}$ as the identity matrix, and ${\bf h}=(1,0.5)\in{\cal D}_{{\bf Z}}$. Assume for the purpose of contradiction that $Z_{1}\sim VG^{1}(b,\mu,\Sigma)$ under $\mathbb{Q}_{{\bf h}}$ for some $(b,\mu,\Sigma)$, which can be obtained by matching the first 3 central moments with those obtained from (4.8). Then $\nu:=1/b$ can be found in closed form by solving a quadratic equation, giving $\nu=1.03788$ or $3.30906$, so that $\operatorname{Kurt}_{\mathbb{Q}_{{\bf h}}}(Z_{1})=6.11364$ or $12.92718$. On the other hand, (4.8) implies $\operatorname{Kurt}_{\mathbb{Q}_{{\bf h}}}(Z_{1})=9/2$, a contradiction. Thus, since a WVAG process must have VG marginal components, this shows the class of WVAG processes is not closed under the Esscher transform.
@@ 4.3 Analytical formula for cgf of the log return
and ${\bf X}(t)=\boldsymbol{\Lambda}(t)-\log{\bf S}(t)$. Using fact that ${\bf Z}^{*}(t,T)$ is independent of ${\cal F}_{t}$, and (2.10), the conditional distribution of ${\bf R}(t,T){\,|\,}{\cal F}_{t}$ is
${\bf R}(t,T){\,|\,}{\cal F}_{t}$ =$\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\Lambda}(T)+{\bf X}(t)e^{-\lambda\tau}-{\bf Y}(t)+{\bf Z}^{*}e^{-\lambda\tau}$ (4.9)
for some random vector ${\bf Z}^{*}\stackrel{{\scriptstyle D}}{{=}}{\bf Z}^{*}(\tau)$ independent of ${\cal F}_{t}$. Let $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}):=\kappa_{{\bf R}(t,T){\,|\,}{\cal F}_{t}}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ denote the cgf of this conditional distribution.
@@
$\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(T)}(\theta)=(e^{\lambda T}-1)\eta_{k}\theta+\sum_{j=1}^{m}\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(T)}(\theta),$
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(T)}$ is evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.
@@
$\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(T)}(\theta)=(e^{\lambda T}-1)\eta_{k}\theta+\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(T)}(\theta)+\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(T)}(\theta),$
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(T)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(T)}$ are evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.
@@
Note that following all the substitutions, all the cgfs $\kappa^{\mathbb{Q}_{\bf h}}_{{\bf R}(t,T)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}$ presented in this section have analytical formulas in terms of the dilogrothim function.
@@ 4.4.1 Forward prices
Suppose that ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ or ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$, and assume ${\bf h},{\bf e}_{k}\in{\cal D}_{{\bf Z}}$. The forward price of $S_{k}$ at time $t$ with maturity time $T$ is
$F_{t}$ $=S_{t}\exp({\kappa_{R_{k}(t,T)}^{\mathbb{Q}_{\bf h}}(1)}),\quad t\in[0,T],$ (4.13)
@@
$F_{t}=\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(T){\,|\,}{\cal F}_{t}]=S_{t}\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[e^{R_{k}(t,T)}{\,|\,}{\cal F}_{t}].$
@@
Figure 1: Forward price curve $\tau\mapsto F_{8}$ on $S_{1}$, where the market price of risk is ${\bf h}=\gamma(-0.1,-0.03)$, $\gamma=-1,1,10$. All parameters are otherwise taken from Section 5.1 but $b_{21}=b_{31}=0$ for all curves except the one with cyclical component.
@@ 4.4.2 Call options
$c_{t}=C\int_{0}^{\infty}\operatorname{Re}\mathopen{}\left(\frac{S_{k}(t)^{{\rm i}v}K^{-{\rm i}v}\exp({\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}}(\epsilon+1+{\rm i}v))}{\epsilon^{2}+\epsilon-v^{2}+{\rm i}(2\epsilon+1)v}\right)\mathclose{}{\rm d}v,\quad t\in[0,T],$ (4.14)
@@
and $\epsilon>0$ is the dampening parameter ($\alpha$ in CM99 ). Here, the cgf $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ is obtained by Proposition 4.7 or Corollary 4.8 and its finiteness in the integrand is provided by $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ As usual, ${\bf h}\in{\cal D}_{{\bf Z}}$ is necessary for $\mathbb{Q}_{{\bf h}}$ to exist. Provided these conditions hold, $\epsilon$ has no effect on the call price. We make further comments on the choice of $\epsilon$ in Section 5.1.
@@
Note that $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ is equivalent to $\mathbb{E}_{\mathbb{Q}_{\bf h}}[S_{k}^{\epsilon+1}]<\infty$, which is the necessary and sufficient condition for the Carr-Madan formula to be finite. For example, if the BDLP is ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ then by combining bound for $\epsilon$ in (CM99, , Section 5), Corollary 4.8, and (2.5), this is equivalent to
$1\displaystyle 0<\epsilon<\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}-$ 1
by noting that $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf V}_{0}}$ and $(\epsilon+1){\bf e}_{k}\in{\cal D}_{V_{k}}$ turn out to be the same condition, and recalling the domain defined using the parameters under $\mathbb{P}$.  $\Box$
@@ 4.4.3 Spread options
Consider a spread option on the biviarate energy prices $(S_{k},S_{l})$, $k,l=1,\dots,n$, $k\neq l$, with strike price $K$ and maturity time $T$. By Remark 4.4, if $n\geq 3$, this is a more general setting with the possibility that the components ${\bf h}$ other than the $k$th and $l$th component affect the law of $(S_{k},S_{l})$ under $\mathbb{Q}_{{\bf h}}$.
@@
$t\in[0,T]$, expressed as a function of the current energy price $(S_{k},S_{l})(t)$ and strike price $K$, which is possible as the LDOUP is Markov and the option is path independent. Suppose also that $-\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$, then the Hurd-Zhou formula HZ10 states that the price with strike 1 is
$\begin{split}&f_{t}((S_{k},S_{l})(t),1)=\\ &\quad\quad\frac{e^{-rT}}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\exp(\kappa_{(R_{k},R_{l})(t,T)}^{\mathbb{Q}_{{\bf h}}}(-\boldsymbol{\epsilon}+{\rm i}\boldsymbol{\theta}))e^{\langle-\boldsymbol{\epsilon}+{\rm i}\boldsymbol{\theta},{\bf Y}(t)\rangle}\hat{P}(\boldsymbol{\theta}+{\rm i}\boldsymbol{\epsilon})\,{\rm d}\boldsymbol{\theta},\end{split}$ (4.16)
@@
$\hat{P}(\boldsymbol{\theta})=\frac{\Gamma({\rm i}(\theta_{k}+\theta_{l})-1)\Gamma(-{\rm i}\theta_{l})}{\Gamma({\rm i}\theta_{k}+1)},\quad\boldsymbol{\theta}=(\theta_{k},\theta_{l}),$
$\Gamma$ is the gamma function, and $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})$ are dampening parameters that satisfy $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$. Here, $\kappa^{\mathbb{Q}_{\bf h}}_{(R_{k},R_{l})(t,T)}(\boldsymbol{\theta})$ can be obtained by Proposition 4.5 or Corollary 4.6 by taking the $k$th and $l$th marginal component if $n\geq 3$ similar to Proposition 4.7, and its finiteness in the integrand is provided by $-\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$. As before, provided these conditions hold, $\boldsymbol{\epsilon}$ has no effect on the spread option price. Then the spread option price for a general strike price $K$ is
$f_{t}(S_{k}(t),S_{l}(t),K)=Kf_{t}\mathopen{}\left(\frac{(S_{k},S_{l})(t)}{K},1\right)\mathclose{}.$ (4.17)

### v2

@@ 4.1 Energy market model
$k=1,\dots,n$, where $p$ is the period, and $b_{0k},b_{1k},b_{2k},b_{3k}\in\mathbb{R}$. Let
${\bf X}\sim OU\text{-}VGC^{n}_{m}(\lambda,{\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$
with initial value ${\bf X}_{0}$ in its stationary distribution. Thus, the BDLP is ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, and we work within this more general setting throughout, which includes ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a subclass. In Section 5.1, we make an additional assumption on $\boldsymbol{\eta}$, which is not needed now.
@@ 4.2 Esscher transform
It is immediate that $\mathbb{Q}_{{\bf h}}$ exists if and only if ${\bf h}\in{\cal D}_{\bf Z}$ and it is equivalent to $\mathbb{P}$. The Esscher transform is structure preserving in the sense that $\bf Z$} is also a Lévy process under $\mathbb{Q}_{\bf h}$.
The no arbitrage condition that the discounted traded asset prices are $\mathbb{Q}_{{\bf h}}$-martingales trivially holds for the risk-free asset, which is the only traded asset, so that the Esssher transform provides up to infinitely many EMMs $\mathbb{Q}_{{\bf h}}$ indexed by ${\bf h}\in{\cal D}_{\bf Z}$, all consistent with no arbitrage. The market chooses a MPR ${\bf h}\in{\cal D}_{{\bf Z}}$ based on risk preferences, which we infer by calibrating the model to observed forward or call prices.
@@
Without loss of generality, suppose $\boldsymbol{\eta}={\bf 0}$. Let ${\bf h}\in{\cal D}_{{\bf Z}}$. Combining (shir, , Section VII.2c, Theorem 1), Kac’s theorem, and that ${\bf V}_{1},\dots,{\bf V}_{m}$ are independent under $\mathbb{P}$, it follows that they are also independent under $\mathbb{Q}_{{\bf h}}$. Then applying Lemma 4.1 to each ${\bf V}_{1},\dots,{\bf V}_{m}$ yields its law under under $\mathbb{Q}_{{\bf h}}$.
Note that (4.4) holds since $\mathbb{E}[e^{\langle\boldsymbol{\theta},{\bf V}_{1}+\dots+{\bf V}_{m}\rangle}]=\prod_{j=1}^{m}\mathbb{E}[e^{\langle\boldsymbol{\theta},{\bf V}_{j}\rangle}]$ with the LHS finite if and only if each term in the product is, otherwise both sides are infinite. ∎
@@
One way to see this is to take the example $n=2$, $a=0.5$, $\boldsymbol{\alpha}=(1,1)$, $\boldsymbol{\mu}=\boldsymbol{\eta}={\bf 0}$, $\boldsymbol{\Sigma}$} as the identity matrix, and ${\bf h}=(1,0.5)\in{\cal D}_{{\bf Z}}$. Assume for the purpose of contradiction that $Z_{1}\sim VG^{1}(b,\mu,\Sigma)$ under $\mathbb{Q}_{{\bf h}}$ for some $(b,\mu,\Sigma)$, which can be obtained by matching the first 3 central moments with those obtained from (4.8). Then $\nu:=1/b$ can be found in closed form by solving a quadratic equation, giving $\nu=1.03788$ or $3.30906$, where only the former satisfies the parameter constraints including $\Sigma>0$, so that $\operatorname{Kurt}_{\mathbb{Q}_{{\bf h}}}(Z_{1})=8.97664$. On the other hand, (4.8) implies $\operatorname{Kurt}_{\mathbb{Q}_{{\bf h}}}(Z_{1})=9.07618$, a contradiction. Thus, since a WVAG process must have VG marginal components, the class of WVAG processes is not closed under the Esscher transform.
@@ 4.3 Analytical formula for cgf of the log return
and ${\bf X}(t)=\boldsymbol{\Lambda}(t)-\log{\bf S}(t)$. Using fact that ${\bf Z}^{*}(t,T)$ is independent of ${\cal F}_{t}$, and (2.10), the conditional distribution ${\bf R}(t,T){\,|\,}{\cal F}_{t}$ satisfies
${\bf R}(t,T){\,|\,}{\cal F}_{t}$ $\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\Lambda}(T)+{\bf X}(t)e^{-\lambda\tau}-{\bf Y}(t)+{\bf Z}^{*}e^{-\lambda\tau}$ (4.9)
for a random vector ${\bf Z}^{*}\stackrel{{\scriptstyle D}}{{=}}{\bf Z}^{*}(\tau)$ independent of ${\cal F}_{t}$. Let $\kappa_{{\bf R}(t,T)}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta}):=\kappa_{{\bf R}(t,T){\,|\,}{\cal F}_{t}}^{\mathbb{Q}_{\bf h}}(\boldsymbol{\theta})$ denote the cgf of this conditional distribution.
@@
$\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(\tau)}(\theta)=(e^{\lambda\tau}-1)\eta_{k}\theta+\sum_{j=1}^{m}\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(\tau)}(\theta),$
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{jk}^{*}(\tau)}$ is evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.
@@
$\kappa^{\mathbb{Q}_{\bf h}}_{Z_{k}^{*}(\tau)}(\theta)=(e^{\lambda\tau}-1)\eta_{k}\theta+\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(\tau)}(\theta)+\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(\tau)}(\theta),$
and $\kappa^{\mathbb{Q}_{\bf h}}_{V_{0k}^{*}(\tau)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{V_{k}^{*}(\tau)}$ are evaluated by (3.7). Furthermore, (4.12) holds for $\theta\in\mathbb{C}$ such that $\operatorname{Re}(\theta){\bf e}_{k}\in{\cal D}_{{\bf Z}}$.
@@
Remark 4.9.
The condition $\theta{\bf e}_{k}\in{\cal D}_{\bf Z}$ in Corollary 4.8 is equivalent to
$-\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}<\theta<\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}$ (4.13)
by recalling ${\cal D}_{{\bf Z}}$ is defined using the parameters under $\mathbb{P}$ in (2.5).   $\Box$
After all the substitutions, all the cgfs $\kappa^{\mathbb{Q}_{\bf h}}_{{\bf R}(t,T)}$, $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}$ presented in this section have analytical formulas in terms of the dilogarithm function.
@@ 4.4.1 Forward prices
Suppose that ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ or ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$, and assume ${\bf h},{\bf e}_{k}\in{\cal D}_{{\bf Z}}$. The forward price of $S_{k}$ at time $t$ with maturity time $T$ is
$F_{t}$ $=S_{k}(t)\exp({\kappa_{R_{k}(t,T)}^{\mathbb{Q}_{\bf h}}(1)}),\quad t\in[0,T],$ (4.14)
@@
$F_{t}=\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(T){\,|\,}{\cal F}_{t}]=S_{k}(t)\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[e^{R_{k}(t,T)}{\,|\,}{\cal F}_{t}].$
@@
Figure 1: Forward price curve $\tau\mapsto F_{8}$ on $S_{1}$, where the MPR is ${\bf h}=\gamma(-0.1,-0.03)$, $\gamma=-1,1,10$. All parameters are otherwise taken from Section 5.1 but $b_{21}=b_{31}=0$ for all curves except the one with cyclical component.
@@ 4.4.2 Call options
$c_{t}=C\int_{0}^{\infty}\operatorname{Re}\mathopen{}\left(\frac{S_{k}(t)^{{\rm i}\theta}K^{-{\rm i}\theta}\exp({\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}}(\epsilon+1+{\rm i}\theta))}{\epsilon^{2}+\epsilon-\theta^{2}+{\rm i}(2\epsilon+1)\theta}\right)\mathclose{}{\rm d}\theta,\quad t\in[0,T],$ (4.15)
@@
and $\epsilon>0$ is the damping parameter ($\alpha$ in CM99 ). Here, the cgf $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ is obtained by Proposition 4.7 or Corollary 4.8 and its finiteness in the integrand is provided by $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ . As usual, ${\bf h}\in{\cal D}_{{\bf Z}}$ is necessary for $\mathbb{Q}_{{\bf h}}$ to exist. Provided these conditions hold, $\epsilon$ has no effect on the call price. We make further comments on the choice of $\epsilon$ in Section 5.1.
@@
Note that $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ is equivalent to $\mathbb{E}_{\mathbb{Q}_{\bf h}}[S_{k}^{\epsilon+1}]<\infty$, which is the necessary and sufficient condition for the Carr-Madan formula to be finite. For example, if the BDLP is ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, then this is equivalent to
$0<\epsilon<\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}-1,$
which follows from (4.13), or alternatively, (CM99, , Section 5).  $\Box$
@@ 4.4.3 Spread options
Consider a spread option on the biviarate energy prices $(S_{k},S_{l})$, $1,\dots,n, klk\neq l$, with strike price $K$ and maturity time $T$. By Remark 4.4, if $n\geq 3$, this is a more general setting with the possibility that the components $\bf h$} other than the $k$th and $l$th component affect the law of $(S_{k},S_{l})$ under $\mathbb{Q}_{{\bf h}}$.
@@
$t\in[0,T]$, expressed as a function of the current energy price $(S_{k},S_{l})(t)$ and strike price $K$, which is possible as the LDOUP is Markov and the option is path independent. Suppose also that $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$, then the Hurd-Zhou formula HZ10 gives the price with strike 1 as
$\begin{split}&f_{t}((S_{k},S_{l})(t),1)=\\ &\quad\quad\frac{e^{-rT}}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\exp(\kappa_{(R_{k},R_{l})(t,T)}^{\mathbb{Q}_{{\bf h}}}(-\boldsymbol{\epsilon}+{\rm i}\boldsymbol{\theta}))e^{\langle-\boldsymbol{\epsilon}+{\rm i}\boldsymbol{\theta},(Y_{k},Y_{l})(t)\rangle}\hat{P}(\boldsymbol{\theta}+{\rm i}\boldsymbol{\epsilon})\,{\rm d}\boldsymbol{\theta},\end{split}$ (4.17)
@@
$\hat{P}(\boldsymbol{\theta})=\frac{\Gamma(-1+{\rm i}(\theta_{k}+\theta_{l}))\Gamma(-{\rm i}\theta_{l})}{\Gamma(1+{\rm i}\theta_{k})},\quad\boldsymbol{\theta}=(\theta_{k},\theta_{l}),$
$\Gamma$ is the gamma function, and $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})$ are damping parameters that satisfy $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$. Here, $\kappa^{\mathbb{Q}_{\bf h}}_{(R_{k},R_{l})(t,T)}(\boldsymbol{\theta})$ can be obtained by Proposition 4.5 or Corollary 4.6 by taking the $k$th and $l$th marginal component if $n\geq 3$, similar to Proposition 4.7, and its finiteness in the integrand is provided by $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$. As before, provided these conditions hold, $\boldsymbol{\epsilon}$ has no effect on the spread option price. Then the spread option price for a general strike price $K$ is
$f_{t}((S_{k},S_{l})(t),K)=Kf_{t}\mathopen{}\left(\frac{(S_{k},S_{l})(t)}{K},1\right)\mathclose{}.$ (4.18)